A stone thrown upwards, has its equation of motion s=490t-4.9t^2 where 's' in metres and t in seconds respectively. what is the maximum height reached by it?

Grade:11

1 Answers

Shivam Taneja
6 years ago
Dear Ameya,
see s=ut-0.5gt^2
comparing this we get u=490
thus for max hieght,
v^2-u^2=2as
v=0 at highest point and a=-9.8
s=490*490/2/9.8
 =490*25
 =12250